# Optimal or Greedy Decision Trees? Revisiting their Objectives, Tuning, and Performance

### v1

@@ 3.2 Analysis
To further understand the differences between all the objectives listed above, we plot them and analyze the differences. The left of Fig. 2 shows the values of the objectives presented above when the leaf node size is fixed but the number of misclassifications increases. The accuracy is a straight line since every misclassification is counted equally. Both the pessimistic binomial score and the minimum error follow the accuracy tightly, with only a small additional cost for higher misclassifications. All other objectives follow roughly the same pattern: the first misclassifications in a node are penalized most and the additional penalty for extra misclassifications decreases.
@@
Intuitively, it seems odd that the objective increases for larger nodes with a fixed error. Table 1 shows some examples of relative objective values that are unexpected. For example, according to entropy, it is better to have two nodes of size 4, with two misclassifications in the first node and zero in the second, than one node of size 8 with one misclassification. Entropy strongly values a pure node, even if this means a higher misclassification rate. Other objectives, such as MDL, value two nodes of size four and two with two misclassifications in the first and none in the second, more than one node of size six with also two misclassifications. Again, a small pure node is valued, even though the node of size four with two misclassifications has a high probability of being misclassified.
Table 1: Pure nodes are overvalued, resulting in splits with pure nodes (e.g., $[n,e]=[4,0]$) and nodes that are labeled randomly ($[4,2]$), rather than keeping one node with the same misclassifications ($[8,2]$) or even less in case of entropy ($[8,1]$).
@@
Gini impurity $f(8,2)<f(4,2)+f(4,0)$ $f(8,2)=3.000$, $f(4,2)+f(4,0)=2.000$
Entropy $f(8,1)<f(4,2)+f(4,0)$ $f(8,1)=2.174$, $f(4,2)+f(4,0)=2.000$
MDL (Quinlan) $f(6,2)<f(4,2)+f(2,0)$ $f(6,2)=4.708$, $f(4,2)+f(2,0)=4.377$
MDL (Mehta) $f(6,2)<f(4,2)+f(2,0)$ $f(6,2)=5.513$, $f(4,2)+f(2,0)=5.409$
Bayes $f(6,2)<f(4,2)+f(2,0)$ $f(6,2)=4.379$, $f(4,2)+f(2,0)=4.321$
@@ 3.3 Novel Non-Concave Objectives
The odd behavior of the greedy criteria in Fig. 2 and Table 1 is a result of their strict concavity. (Strict) concavity is not a requirement for ODTs because ODTs do not consider splitting criteria and can search beyond a non-improving split. Therefore we here introduce four non-concave objectives that have not previously been used in decision tree learning.
Error Gini:
Instead of weighing the Gini impurity by the leaf node size, we propose Error Gini which multiplies the Gini impurity by the number of errors in a leaf node. This ensures that the objective value does not increase when the misclassifications are fixed but the size of the leaf node increases. Again, let $p_{0}=\frac{e}{n}$ and $p_{1}=\frac{n-e}{n}$, then the objective value is
$f(n,e)=e(1-p_{0}^{2}-p_{1}^{2})\,.$
Error entropy:
Similarly, we provide Error entropy which gives the entropy multiplied by the number of errors, instead of weighted by the size of the leaf node:
$f(n,e)=-e(p_{0}\log_{2}p_{0}+p_{1}\log_{2}p_{1})\,.$
@@
Noel et al. (2023) propose two new objective functions for optimizing neural networks, both of which are non-concave. The first new objective function they propose is the M-loss, here rewritten in terms of $n$ and $e$:
@@
$f(n,e)=n\left(\frac{1}{\sqrt{1-\left(\frac{e}{n}\right)^{2}}}-1\right)\,.$
Fig. 3 shows the values these new functions take. For easy comparison, accuracy and Gini impurity are also included in the plot. In contrast to the strictly concave functions, the left side of Fig. 3 shows how the first misclassifications in a leaf node are penalized less, whereas nodes with a (close to) balanced class distribution are heavily penalized. The right side shows that increasing the leaf node size while keeping the number of misclassifications constant, decreases the penalization. Therefore, these objectives obtain the desired property to penalize nodes with a close-to-equal class distribution more strongly.
@@ 3.4 Experiments
    Which objectives are well suited for small data sets?
@@ 3.4.1 Experiment Setup
We empirically compare the objectives on a large benchmark set from OpenML (Vanschoren et al., 2013; Feurer et al., 2021). [^2] The code will be made public on paper publication. For the sake of scalability, we selected all binary classification data sets with 50 or fewer features, of which ten or fewer numeric features and no large text features, with no missing values, with at most 100,000 instances, and at least 40 instances. We take the most recent version of the data set and omit duplicates or data sets that only differ in the random seed, resulting in 165 data sets. We split each data set into five folds, creating five train and test pairs each consisting of four and one fold respectively. We list all data sets used in Appendix A.
We implemented all the ODT objectives in the ODT method STreeD (van der Linden et al., 2023) because of its scalability and flexibility in supporting new objectives. [^3] https://github.com/algtudelft/pystreed STreeD is a DP approach that requires binary features. Therefore we binarize the numeric training data with thresholds on the ten quantiles, and the categorical data with one-hot encoding (with at most ten categories). The test data is binarized in the same way. We experimented with other binarization approaches but noted no significant impact. See Section 5.2 for the impact of binarization on CART.
We train ODTs up to a depth of three and four while tuning both the tree depth and the number of branching nodes using five-fold cross-validation. A depth-three tree has at most eight leaf nodes and seven branching nodes and a depth-four tree has at most 16 leaf nodes and 15 branching nodes.
@@
Figure 4: Comparing ODT objectives for (a) $\operatorname{max-depth}=3$ and (b) $\operatorname{max-depth}=4$. Orange (blue) indicates (non-)concave. The average rank, accuracy, and number of leaf nodes over all data sets and folds are shown.
Figure 5: Nemenyi critical distance rank test for ODTs of maximum depth four. The average rank per objective is plotted and objectives with a rank difference smaller than the critical distance (CD) at p-value $0.05$ are grouped by a black bar.
3.4.2 Optimal Decision Tree Results
Figs. 4 and 4 show the average rank, test accuracy, and the number of leaf nodes per objective for trees of depth three and four. On average, the best ranking objectives with the best test accuracy are accuracy and its slight variations Minimum Error and the Binomial Pessimistic error, and our novel non-concave functions M-loss, error entropy, error Gini, and L-loss. Maximizing the training accuracy is the second and highest-ranked objective for maximum depths three and four respectively. However, the highest average test accuracy for both depth limits is obtained by minimizing the new objective M-loss. For a maximum depth of four, the pessimistic binomial objective achieves a similar average test accuracy with a lower number of leaf nodes than the accuracy objective.
However, the results are close. Fig. 5 shows the results of a Nemenyi critical distance rank test to test the significance (Demšar, 2006). This test computes the critical distance (CD) between the average ranks of two methods to be statistically significant. Fig. 5 shows that all the non-concave objectives and Gini are not significantly different for depth four. The non-concave objectives are significantly better than the concave objectives (except Gini).
3.4.3 Small Data Sets
We noted that the differences among objectives were larger for smaller data sets. These data sets are particularly hard because overfitting is likely with such limited information. Therefore Fig. 6 shows the results on the 70 data sets with less than 200 instances.
Figure 6: Comparing ODT objectives for $\operatorname{max-depth}=4$ on 70 small data sets ($n<200$).
The new non-concave objectives perform significantly better than accuracy, with the test accuracy of M-loss on average exceeding that of the accuracy objective by $0.7\%$. Therefore, with sufficient data, ODTs obtain the best performance by maximizing training accuracy but for smaller data sets, other objectives, such as the M-loss can yield better results.
3.4.4 Greedy Heuristics Results
Figure 7: Comparing greedy objectives for $\operatorname{max-depth}=4$. The strictly concave objectives outperform the non-concave objectives.
For comparison, we also train greedy trees with the objectives listed above. Fig. 7 shows the resulting out-of-sample performance. In comparison with the ODTs, the ranking of the objectives is almost reversed, which is in line with our analysis in Section 3.2. A Nemenyi critical distance rank test shows that all strictly concave objectives are significantly better than all non-concave objectives for TDI heuristics. The smaller tree sizes of the non-concave objectives show that the greedy heuristic gets stuck early with these objectives and finds no improving splits. Optimizing Gini impurity yields 1.2% higher out-of-sample accuracy than directly optimizing the in-sample accuracy. These results confirm that the traditional Gini and entropy are among the top choices.
@@ 3.5 Discussion
Previous work has extensively compared greedy splitting criteria, which we extend to optimal decision trees. Our results show that optimizing accuracy directly is a good choice for ODTs, specifically when the training data is large because, with sufficient training data, the training accuracy approximates the test accuracy. This shows that the strict concavity of objectives such as Gini impurity and entropy, is not an inherently necessary or desirable property, but a limitation imposed by the greedy TDI approach. Optimal and greedy training procedures respond differently to the objectives and therefore best practices of one approach do not necessarily translate to the other.
We obtained better results for small data sets by using one of our new non-concave objectives, such as M-loss or error entropy. These objectives strongly penalize leaf nodes with a close to equal class distribution and are therefore more robust when there is little data available. This does not apply to greedy heuristics which show an almost opposite ranking of objectives in comparison to ODTs. This opens the question of how the performance of non-concave objectives for small data sets could be exploited in greedy heuristics.
4 Tuning the Tree Complexity of Optimal Decision Trees
Most decision tree learning approaches make a trade-off between training accuracy and model complexity to prevent overfitting. For ODTs, several complexity tuning methods have been used but without an in-depth empirical comparison. Therefore this section analyses the effect of complexity tuning methods for optimal decision trees, starting with an overview of complexity tuning methods. We find that optimal decision trees perform significantly better with tuning than without, but that existing tuning techniques perform similarly.
@@ 4.1 Tuning Approaches
The following complexity tuning approaches are currently used in ODT methods:
Depth:
A common metric for tree complexity is its depth. DL8.5 (Aglin et al., 2020a, b), for example, tunes the depth of the tree. A possible disadvantage is that the number of nodes increases exponentially with respect to the maximum depth, thus providing only a coarse control of the tree size. In our experiments below, we tune the depth parameter $d\in\{0,...,\operatorname{max-depth}\}$, with a $\operatorname{max-depth}$ of either four or five.
Size:
The number of nodes can be tuned directly as a hard constraint. Since binary trees always have one more leaf node than branching nodes, the total number of nodes can also be counted by the number of branching or leaf nodes alone. This approach is taken by Demirović et al. (2022) and van der Linden et al. (2023) who directly tune both the maximum depth and the number of branching nodes. In our experiments below, we tune the depth and the number of nodes such that the depth $d\in\{0,...,\operatorname{max-depth}\}$ and the number of branching nodes $n\in\{d,...,2^{d}-1\}$.
Complexity cost:
The most common approach in greedy trees is to penalize the cost of adding a node by a factor $\lambda$. This approach is used in most MIP methods (Bertsimas and Dunn, 2017) and also in the optimal sparse approaches by Lin et al. (2020) and Chaouki et al. (2024). Complexity-cost tuning minimizes $\lambda(|L|-1)+\sum_{(n,e)\in L}f(n,e)$, with $L$ the set of leaves and $\lambda$ the complexity-cost parameter.
As similarly done in (Lin et al., 2020), we tune the complexity-cost parameter $\lambda\in\{0.0001$, $0.0002$, $0.0003$, $0.0004$, $0.0005$, $0.0006$, $0.0007$, $0.0008$, $0.0009$, $0.001$, $0.002,0.003$, $0.004$, $0.005$, $0.006$, $0.007$, $0.008$, $0.009$, $0.01$, $0.02$, $0.03$, $0.04,0.05$, $0.06$, $0.07$, $0.08$, $0.09$, $0.1$, 0.2}0.2\}0.2 } in our experiments below.
Minimum Support:
The minimum support or minimum leaf node size is a hard constraint on the minimum number of instances that should end up in a leaf node. Closely related is the minimum split size: the minimum number of instances in a node to allow for a split. Minimum support is more common in the data mining literature, e.g., Nijssen and Fromont (2007). It is also introduced as a soft constraint (Vilas Boas et al., 2021).
In our experiments below, we tune the minimum leaf node size $m$ with possible values for $m$ expressed as a percentage $m^{\prime}$ of the data set size $n$ with $m^{\prime}\in\{0.001$, $0.002$, $0.003$, $0.004$, $0.005$, $0.0075$, $0.01$, $0.015$, $0.02$, $0.03$, $0.04$, $0.05$, $0.06$, $0.07$, $0.08$, $0.09$, $0.1$, $0.15$, $0.2$, $0.25$, $0.3$, $0.35$, $0.4$, $0.45$, 0.5}0.5\}0.5 }. Rounding $m^{\prime}n$ to an integer may yield non-unique numbers, so we select the smallest ten unique integer possible values for $m$.
In our experiments, we leave out selecting a tree structure from a fixed set of possible structures (Günlük et al., 2021), because this approach was chosen specifically to deal with the scalability limitations of MIP.
4.2 New Tuning Approaches
On top of the existing ODT tuning approaches, we evaluate three new approaches:
@@ 4.2 New Tuning Approaches
G-Test threshold:
The G-Test threshold approach measures the difference between the two branches of a split and only allows splits that result in splits with a statistically significant difference. The G-Test is a likelihood ratio test that works analogously to the $\chi^{2}$ test used in CHAID (Kass, 1980) and ID3 (Quinlan, 1986). In our experiments below, we tune the significance level $\alpha\in\{0.000001$, $0.00001$, $0.0001$, $0.001$, $0.005$, $0.01$, $0.05$, $0.1$, $0.2$, $0.5$, 1.0}1.0\}1.0 }.

### v2

@@ 3.2 Analysis of ODT Objectives
The left of Fig. 2 shows the values of these objectives when the leaf node size is fixed but the number of misclassifications increases. The accuracy is a straight line since every misclassification is counted equally. Both the pessimistic binomial score and the minimum error follow the accuracy tightly, with only a small additional cost for higher misclassifications. All other objectives follow roughly the same pattern: the first misclassifications in a node are penalized most and the additional penalty for extra misclassifications decreases.
@@
It is counter-intuitive that the objective increases for larger nodes with a fixed error. Table 1 shows some examples of relative objective values that are unexpected. For example, according to the entropy criterion, it is better to have two nodes of size 4, with two misclassifications in the first node and zero in the second, than one node of size 8 with one misclassification. Entropy strongly values a pure node, even if this means a higher misclassification rate. Other objectives, such as MDL, value two nodes of size four and two with two misclassifications in the first and none in the second, more than one node of size six with also two misclassifications. Again, a small pure node is valued, even though the node of size four with two misclassifications has a high probability of being misclassified, for example, in the presence of class noise.
Table 1: Pure nodes are overvalued, resulting in splits with pure nodes (e.g., $(n,e)=(4,0)$) and nodes that are labeled randomly $(4,2)$, rather than keeping one node with the same misclassifications $(8,2)$ or even less in case of entropy $(8,1)$.
@@
Gini impurity $f(8,2)\leq f(4,2)+f(4,0)$ $f(8,2)=3.000$, $f(4,2)+f(4,0)=2.000$
Entropy $f(8,1)\leq f(4,2)+f(4,0)$ $f(8,1)=2.174$, $f(4,2)+f(4,0)=2.000$
MDL (Quinlan) $f(6,2)\leq f(4,2)+f(2,0)$ $f(6,2)=4.708$, $f(4,2)+f(2,0)=4.377$
MDL (Mehta) $f(6,2)\leq f(4,2)+f(2,0)$ $f(6,2)=5.513$, $f(4,2)+f(2,0)=5.409$
Bayes $f(6,2)\leq f(4,2)+f(2,0)$ $f(6,2)=4.379$, $f(4,2)+f(2,0)=4.321$
@@ 3.3 Novel Non-Concave Objectives
The odd behavior of the greedy criteria in Fig. 2 and Table 1 is a result of their strict concavity. (Strict) concavity is not a requirement for ODTs because ODTs do not consider splitting criteria and can search beyond a non-improving split. Therefore, we here introduce two non-concave objectives by Noel et al. (2023) that have not previously been used in decision tree learning.
@@
The first is the M-loss, here rewritten in terms of $n$ and $e$:
@@
$f(n,e)=n\Bigg{(}\frac{1}{\sqrt{1-\left(\frac{e}{n}\right)^{2}}}-1\Bigg{)}\,.$
Fig. 3 shows the values these new functions take. For easy comparison, accuracy and Gini impurity are also included in the plot. In contrast to the strictly concave functions, the left side of Fig. 3 shows how the first misclassifications in a leaf node are penalized less, whereas nodes with a (close to) balanced class distribution are heavily penalized. The right side shows that increasing the leaf node size while keeping the number of misclassifications constant, decreases the penalization. Therefore, we hypothesize that these objectives obtain the desired property to penalize nodes with a close-to-equal class distribution more strongly.
@@ 3.4 Experiments
    How do different objectives respond to noise and data set size?
@@ 3.4.1 Experiment Setup
We empirically compare the objectives on a large benchmark set from OpenML (Vanschoren et al., 2013; Feurer et al., 2021). [^2] The code will be made public on paper publication. For the sake of scalability, we selected all binary classification data sets with 50 or fewer features, of which eight or fewer numeric features and no large text features, with no missing values, with at most 100,000 instances, and at least 40 instances. We take the most recent version of the data set and omit duplicates or data sets that only differ in the random seed, resulting in 180 data sets. We split each data set into five folds, creating five train and test pairs each consisting of four and one fold respectively. We list all data sets used in Appendix E and also include a histogram of the data set sizes.
We implemented all the ODT objectives in the ODT method STreeD (van der Linden et al., 2023) because of its scalability and flexibility in supporting new objectives. [^3] https://github.com/algtudelft/pystreed STreeD is a DP approach that requires binary features. Therefore, we binarize the numeric training data with thresholds on the ten quantiles, and the categorical data with one-hot encoding (with at most ten categories). The test data is binarized in the same way. We experimented with other binarization approaches, but noted no significant impact with regard to the analysis presented here. See Section 5.2 for the impact of binarization on CART. See Appendix D.2 for a brief evaluation of the impact of binarization on ODTs.
Additionally, we evaluate on synthetic data sets where we can control the amount of noise. We follow the synthetic data setup from Murthy and Salzberg (1995); Bertsimas and Dunn (2017) and Dunn (2018). We generate $n$ random training instances with $p$ numeric features, uniformly distributed over $[0,1]$. For a given noise strength $f$, we add feature noise by adding noise uniformly drawn from $[-f,f]^{p}$. Again, we binarize the numeric features by threshold predicates on 10 quantiles per numeric feature. We generate a random binary tree on this binarized data of a maximum depth $d$ with at most $2^{d}$ leaf nodes. We choose random splits on the data such that each leaf node contains at least 5 instances. The binary labels of each leaf node are assigned alternately, such that no split leads to two leaf nodes with the same label. After this, we add class noise to a given percentage $c$ of the data by flipping its label. For each training set, we create a corresponding test set without noise of $1000$ instances per leaf node in the generated tree.
While keeping the other values constant, we test with changing the amount of feature noise ($f=0$, $0.2$, $0.4$, $0.6$, $0.8$, $1.0$) and the amount of class noise ($c=0\%$, $10\%$, $20\%$, $30\%$, $40\%$, $50\%$), while also changing the number of instances ($n=50$, $100$, $1000$). We repeat each configuration 1000 times and report averages over these 1000 runs.
We train ODTs up to depth four while tuning the number of branching nodes using cross-validation. [^4] A depth-three tree has at most eight leaf nodes and seven branching nodes and a depth-four tree has at most 16 leaf nodes and 15 branching nodes. In cross-validation, for data sets with up to 100 instances, we use 20 folds; for up to 250 instances, we use ten folds; and otherwise, we use five.
Figure 4: Comparing ODT objectives for $\operatorname{max-depth}=4$. (a) Orange (blue) indicates (non-)concave. The average accuracy and number of leaf nodes over all data sets and folds are shown, sorted by the average rank. (b) Nemenyi critical distance rank test. The average rank per objective is plotted and objectives with a rank difference smaller than the critical distance (CD) at p-value $0.05$ are grouped by a black bar.
@@
The average rank is our main performance metric: for each data set split, we round the test accuracy to one decimal and then rank all methods. If multiple methods have the same accuracy, they are all assigned the average rank. E.g., if two methods have the same best score, they both get rank $1.5$. We then report the mean rank over all data sets.
3.4.2 Optimal Decision Tree Results
Fig. 4 shows the average rank, test accuracy, and the number of leaf nodes per objective for trees of depth four on the OpenML benchmark. On average, the best ranking objectives with the best test accuracy are accuracy and its slight variations minimum error and the binomial pessimistic error, and the novel non-concave functions M-loss and L-loss. Maximizing training accuracy is ranked second, although the difference from the top objective is not significant. The pessimistic binomial objective achieves the best test accuracy with a lower average number of leaf nodes than the accuracy objective.
However, the results are close. Fig. 4b shows the results of a Nemenyi critical distance rank test to test the significance (Demšar, 2006). This test computes the critical distance (CD) between the average ranks of two methods to be statistically significant. Fig. 4b shows that all the non-concave objectives are not significantly different for depth four. All the non-concave objectives are significantly better than all the concave objectives.
The runtimes of optimizing the objectives are close: the objective with the lowest runtime (L-loss) is on average 1.9 times faster (geometric mean) than the slowest objective (Bayes).
In Appendix D.1, we show training accuracy results for a selection of data sets for three objectives (accuracy, Gini impurity, and M-loss) for an increasing number of nodes.
3.4.3 Greedy Heuristics Results
Figure 5: Comparing greedy objectives for $\operatorname{max-depth}=4$. The strictly concave objectives (orange) significantly outperform the non-concave objectives (blue).
For comparison, we also train greedy trees with the objectives listed above. Fig. 5 shows the resulting out-of-sample performance. In comparison with the ODTs, the ranking of the objectives is almost reversed, which is in line with our analysis in Section 3.2, that the TDI approach requires strictly concave objectives. The Nemenyi critical distance rank test in Fig. 5b shows that all strictly concave objectives (except square-root Gini) are significantly better than all non-concave objectives for TDI heuristics. The smaller tree sizes of the non-concave objectives show that the greedy heuristic gets stuck early with these objectives and finds no improving splits. Optimizing Gini impurity yields 1.2% higher out-of-sample accuracy than directly optimizing the in-sample accuracy. These results confirm that the traditional Gini and entropy are among the top choices.
3.4.4 Noisy Synthetic Data
To further investigate the performance of the ODT objectives, we compare their relative performance on synthetic data for varying data set sizes and levels of feature and class noise. We limit the comparison to the top four objectives and include Gini impurity as the top representative of the concave objectives.
Fig. 6 shows that for the smallest data sets ($n=50$) with an increasing level of feature noise, the binomial pessimistic and the minimum error obtain (statistically) significantly better test accuracy ($+1\%$) than the other objectives. Gini impurity performs slightly worse. With more data ($n=100,1000$), the advantage of the binomial pessimistic and minimum error disappears, whereas Gini impurity performs relatively even worse, up to $2\%$.
Fig. 6 shows that the differences for increasing amounts of class noise are smaller. However, Gini impurity still performs worse than the other objectives, especially for larger data sets ($n=1000$) and more class noise. When the class noise reaches $50\%$, half of the labels are flipped and the training labels essentially are completely random. That is why the relative test accuracy for $50\%$ class noise goes back to zero.
@@ 3.4.4 Noisy Synthetic Data
Figure 6: Relative test accuracy of ODT objectives compared to optimizing training accuracy directly. The concave Gini impurity performs significantly worse. For small datasets with feature noise, the binomial pessimistic and minimum error perform significantly better (the error bar represents the 95% confidence interval).
@@ 3.5 Discussion
Previous work has extensively compared greedy splitting criteria, which we extend to optimal decision trees. Our results show that optimizing accuracy directly is a good choice for ODTs, specifically when the number of training samples increases because, with sufficient training data, the training accuracy closely approximates the test accuracy. This shows that the strict concavity of objectives such as Gini impurity and entropy, is not an inherently necessary or desirable property, but a limitation imposed by the greedy TDI approach. Optimal and greedy training procedures respond differently to the objectives and therefore best practices of one approach do not necessarily translate to the other. This opens the question of how the performance of non-strictly-concave objectives for small data sets could be exploited in greedy heuristics.
We hypothesized that non-concave objectives may perform better than accuracy for noisy data. Although we did not measure a significant difference for the new objectives M-loss and L-loss, we found two objectives that can outperform optimizing accuracy directly for small noisy data sets: the pessimistic binomial and minimum error objective. Both of these objectives include a regularizing component, which –in addition to the regularization effect of tuning the tree size– helps to prevent overfitting. This also results in slightly smaller models, while retaining the same average accuracy.
4 Tuning the Complexity of Optimal Decision Trees
Most decision tree learning approaches make a trade-off between training accuracy and model complexity to prevent overfitting. For ODTs, several complexity tuning methods have been used but without an in-depth empirical comparison. Therefore, this section analyses the effect of complexity tuning methods for optimal decision trees, starting with an overview of complexity tuning methods. We find that optimal decision trees perform significantly better with tuning than without, but that existing tuning techniques perform similarly.
@@ 4.1 Tuning Approaches
Currently, ODT approaches mostly tune the following hyperparameters: tree depth (Aglin et al., 2020a), tree size (Demirović et al., 2022), complexity cost (Lin et al., 2020), and the minimum support (Nijssen and Fromont, 2007). In Appendix B, we provide additional information on these approaches. On top of the existing ODT tuning approaches, we evaluate two new approaches: